OpenStudy (anonymous):

find the zeros : f(t)=175t-t^2

OpenStudy (anonymous):

do you know what the difference of two squares is?

OpenStudy (luigi0210):

Hello @MCath Welcome to Openstudy :)

OpenStudy (anonymous):

Helllloooooo 175 is 13.228????

OpenStudy (anonymous):

this is not the difference of two square factor as \[f(t)=t(175-t)\] and then it should be straight forward

OpenStudy (anonymous):

oh ok, just to clarify @satellite73 , when is it DOTS?

OpenStudy (luigi0210):

Yup, now just set each equal to 0 and solve :)

OpenStudy (luigi0210):

@seesawn When it has perfect squares: Like (x^2-16) or (x^2-25) Or (4x^2-81)

OpenStudy (jackellyn):

@seesawn when it is \[a^2-b^2\]

OpenStudy (anonymous):

oh, thanks

OpenStudy (anonymous):

how am i going to solve this

OpenStudy (anonymous):

sorry i have no idea about this one T_T

OpenStudy (jackellyn):

@MCath, after you factor it to t(175-t) t=0 and 175-t=0 Solve 175-t=0 for t

OpenStudy (anonymous):

175 isn't a perfect square how am i going to do with that ???

OpenStudy (jackellyn):

This problem doesn't require using perfect square. You just need to factor it and then solve for the zeros.

OpenStudy (anonymous):

ohhhhhhhh so {0,175} ??

OpenStudy (jackellyn):

Here's an example, \[x^2-74x\] x(x-74) x=0 x=74

OpenStudy (anonymous):

right not to repeat myself, but this is NOT the difference of two squares no idea where that came from

OpenStudy (jackellyn):

@MCath , are you sure your sign is correct for 175? Show me your work

OpenStudy (anonymous):

aahhhhh i really got no idea about this one.. but is this correct.. 175t-t^2 t(t-175) x=0 x=175????

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello anybody there??

OpenStudy (mertsj):

1. Set 175t-t^2 = to 0 2. Factor out t 3. Set each factor equal to 0 4. Solve each equation.

OpenStudy (jackellyn):

@MCath , you're right!

OpenStudy (anonymous):

Yehey!! i have another question... please.. y=36x^2+36x+9 am i going to square all the numbers?? and after after squaring??

OpenStudy (anonymous):

after squaring what will i do

OpenStudy (mertsj):

What are you supposed to do with this problem?

OpenStudy (anonymous):

FIND the zeros

OpenStudy (mertsj):

Factor and set it equal to 0. Set each factor equal to 0. Solve each equation. You will find that there is 1 double root since the given trinomial is a trinomial square.

OpenStudy (anonymous):

ohhhh okay thanks so (6x+3),(6x+3) zeros are {-3,-3} ????? am i right??

OpenStudy (mertsj):

no. Did you set each factor equal to 0 and solve?

OpenStudy (mertsj):

\[6x+3=0\] \[6x=-3\] \[x=\frac{-1}{2}\]

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

so the answer is {-1/2,-1/2} ??

OpenStudy (mertsj):

It is a double root. It would not be listed twice in the solution set.

OpenStudy (mertsj):

The double root indicates a point of tangency to the x axis.

OpenStudy (anonymous):

ohhhh okaaayyy thank you

OpenStudy (mertsj):

yw